OpenStudy (anonymous):

simplify...ILL REWRITE IT 9+9(3-6) TO THE SECOND POWER

OpenStudy (anonymous):

\[9+9(3-6)^{2}\]

OpenStudy (anonymous):

3-6=?

OpenStudy (anonymous):

I rewrote it

OpenStudy (anonymous):

90

OpenStudy (anonymous):

i know, i will help u get the answer step by step so what is 3-6 =

OpenStudy (anonymous):

..........?

OpenStudy (anonymous):

3 ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

now (-3)^2 =?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

you add -3 a 2

OpenStudy (anonymous):

no its -3 to the power 2 so (-3)*(-3)=

OpenStudy (anonymous):

that answer would be -9

OpenStudy (anonymous):

negative * negative = positive :D so +9 what is 9*9=?

OpenStudy (anonymous):

lol man im close

OpenStudy (anonymous):

hehe ok what is 9*9

OpenStudy (anonymous):

+18

OpenStudy (anonymous):

nope, wanna its 9 (times) 9 not 9+9

OpenStudy (anonymous):

81

OpenStudy (anonymous):

9+81=?

OpenStudy (anonymous):

172

OpenStudy (anonymous):

no

OpenStudy (anonymous):

90

OpenStudy (anonymous):

\[9 + 9*(3-6)^{2}\] \[9 + 9*(-3)^{2}\] \[9 + 9*9\] \[9 + 81\] \[90\]

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

no problem, hope u understood

OpenStudy (anonymous):

yes i do now

OpenStudy (anonymous):

that's my ans 4th from the top

OpenStudy (anonymous):

@ajit.shirsat help people get the answer not throw out the answer

OpenStudy (anonymous):

haha....ok ok